OpenStudy (ellecullen):

Homework help asap please: f(x) = 3/x-4 g(x) =(3/x)+4 Find f(g(x)) and g(f(x)), then find the f^-1(x) and check using f(f^-1(x) =x and f^-1(f(x) =x

OpenStudy (ellecullen):

@Mr_Perfection_xD

OpenStudy (ellecullen):

@dan815

OpenStudy (ellecullen):

@ChantySquirrel1129**

OpenStudy (chantysquirrel1129**):

Are you in Connexus?

OpenStudy (ellecullen):

no, I not. I'm simply stumped on this question.

OpenStudy (ellecullen):

I have tried solving this one many times, but its not working out.

OpenStudy (chantysquirrel1129**):

I don't think I could be of much help here...I'm practically useless here :/

OpenStudy (chantysquirrel1129**):

try iambatman c:

OpenStudy (ellecullen):

@iambatman , can you please help me with my homework question

OpenStudy (anonymous):

f(g(x)) means you take the function g(x) and plug it into the x of the function f(x), so for example if I have \[f(x) = 3x^2+5\] and \[g(x) = 9x+2\] and we have to find f(g(x)) we take g(x) and plug it into the x of f(x) so we get \[f(g(x)) = 3(9x+2)^2+5\] and then you simplify etc..

OpenStudy (anonymous):

And the \[f^{-1}(x)\] implies inverse, to do that we can follow these steps: Replace f(x) with y Switch x's and y's, so put x where y is and y where x is (originally). Solve for y Replace y with f^-1(x)

OpenStudy (ellecullen):

the lasttime I tried the problem, I had f(g(x) = 3/((3/x) +4) -4